his Quesion:5 pls A certain group of test subjects had puse rates with a mean of 82 4 beats per minute and a standard deviation of 10 9 beats per minute Would it be unusual for one of the test subjects to have a puibse rate of 542 beats per minde Minimum usual value beats per minute (Type an integer or a decimal) Maximum usual value-beats per minute (Type an ineger or a decimal) s 54 2 beats per minute an unsual pulse rate O A. Yes, because it is betweon the minimum and maximum usual value O B. No, because it is between the minimum and maximum usual valoes O C. No, because it is smaller than the minimum usual value O D. Yes, because it is larger than the maximum usual value E. No, because it is larger than the maximum usual value O F. Yes, because it is smaller than the minimum usual value
Answer #1
minimum usual value =μ-2*σ = 60.60
maximum usual value =μ-2*σ = 104.20

option F is correct